A car initially at rest accelerates at 5.22 M/S to the second for 14.9 seconds how far did it go in this time



1
Expert's answer
2021-11-23T10:32:07-0500

We can find the distance traveled by the car as follows:


"d=\dfrac{1}{2}at^2=\dfrac{1}{2}\times5.22\ \dfrac{m}{s^2}\times(14.9\ s)^2=579\ m."